OpenStudy (gabylovesyou):

Bradley and Kelly are out flying kites at a park one afternoon. A model of Bradley and Kelly's kites are shown below on the coordinate plane as kites BRAD and KELY respectively.

OpenStudy (gabylovesyou):

Which statement is correct about the two kites? They are similar because Segment BR to segment DB is 1:2 and Segment KE to segment YK is 1:2. They are not similar because Segment BR to segment DB is 2:1 and Segment KE to segment YK is 1:2. They are similar because Segment BR to segment DB is 2:1 and Segment KE to segment YK is 2:1. They are not similar because Segment BR to segment DB is 1:2 and Segment KE to segment YK is 2:1.

OpenStudy (gabylovesyou):

@ganeshie8

OpenStudy (gabylovesyou):

@hoblos

OpenStudy (gabylovesyou):

@phi

OpenStudy (phi):

when you write a ratio, you do something like: short / long = short / long for example, in the smaller kite, side BR is short and side BD is long if we knew the length of BR was sqr(5) and the length of BD was 2 sqr(5) we would write ratios: \[ \frac{BR}{BD} = \frac{1 \sqrt{5}}{2\sqrt{5}} \] or, after simplifying \[ \frac{BR}{BD} = \frac{1 }{2} \] notice we do NOT write little / big = big/little (we need the same order big / little on both sides of the = sign)